OpenStudy (anonymous):

Complete the ordered pairs so they are solutions of the equation 5x+2y=40 ( ,0) (0, ) ( ,5)

OpenStudy (anonymous):

just put the given value of x,and y from order pait in the equation to get the required value . when y=0 5x+2(0)=40 5x+0=40 5x=40 x=40/5=8 x=8 so first order pair must be (8,0) can you do the rest ?.